Weston wants to buy a one year membership to a golf course.Rolling Hills Golf course charges a base fee of $200 and an additional $15 per round of golf. A majestic view course changes a base fee of $350 and an additional 10$ per round of golf. which golf course should he choose if he plans to play more than 30 rounds of golf?

Answer by Theo(13342)

the breakeven is 30.
if more than 30, the plan with the cheapest incremental cost will be cheapest.
here's how to analyze.

1.
find the breakeven point.
that occurs when 200 + 15x = 350 + 10x
subtract 200 from both sides and subtract 10x from both sides to get:
5x = 150
solve for x to get x = 30

2.
analyze for 30 and analyze for any number greater than 30.
at x = 30:
200 + 15x = 650
350 + 10x = 650
that's your breakeven
at x = 31
200 + 15x = 665
350 + 10x = 660
the 350 + 10x plan gets cheaper.

graph it to see at a glance.
you can see that the blue graph is below the red graph after the breakeven point.
the blue graph is y = 350 + 10x.